OpenStudy (anonymous):

The position vector of a particle moving in the xy-plane at time t is given by p= (3t^2-4t)i+ (t^2+2t)j. The speed of the particle at t=2 is

OpenStudy (anonymous):

speed = || s'(t) ||

OpenStudy (anonymous):

i got the derivative of (6t-4)i+(3t^2-4t)+(2t=2)j+(t^2+2t) what do i do with the o and j?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

but my answer choices does;t involve any variable.

OpenStudy (anonymous):

plug t = 2 in and you'll get a number

OpenStudy (anonymous):

Is 10 one of your answers?

OpenStudy (anonymous):

@eliassaab yes!! but how??

OpenStudy (anonymous):

\[ p(t) =(3 t^2 - 4 t, t^2 + 2 t)\\ p'(t)=( 6 t-4, 2t +2)\\ p'(2)={8,6}\\ ||p'(2)||= \sqrt{ 8^2 +6^2}=\sqrt {100}=10 \]

OpenStudy (anonymous):

p'(2)=(8,6) = 8 i + 6j

OpenStudy (anonymous):

Where are you?

OpenStudy (anonymous):

it is asking the speed, so won't i just find the velocity and plug in?? you used the distance formula, right?

OpenStudy (anonymous):

The speed is the length of the velocity vector. The length of a i + b j is \[\sqrt{a^2 + b^2} \]

OpenStudy (anonymous):

oh, thanks!!

OpenStudy (anonymous):

YW